The average energy of particles in the early phase of the universe, when it was very
hot, was approximately 10 GeV. Calculate the corresponding temperature.
1
Expert's answer
2019-05-15T10:17:30-0400

The average energy of particles "E=10 GeV = 1.6 \cdot 10^{10} J"

k = Boltzmann’s constant "=1.381 \cdot 10^{−23} J/K"

"E=kT\implies T=\frac{E}{k}"

"T = \frac{1.6\cdot 10^{10}}{1.38\cdot 10^{-23}}=1.15 \cdot 10^{14} K"

Answer "T = 1.15 \cdot 10^{14} K"